Two point changes +q and -q are separated by distance d .A particular electric field line emanating from charge +q at an angle of 60° with horizontal line attains maximum in height h above line AB join charges +q and -q on horizontal plane. IF h = ((√x.)(d))/6, then find the value of x.
1
Expert's answer
2019-05-17T11:18:28-0400



Using Figure we get:


"d \sin 60° = h (1)"

In our case,


"h = \frac {\sqrt {x}}{6} (2)"

We put (2) in (1):



"d \sin 60° = \frac {\sqrt {x}}{6} (3)"

Using (3) we get:

x=27


Answer

27